OpenStudy (anonymous):

How to find the quadratic model for (-2,-14) (0,2) (4,-38)

OpenStudy (anonymous):

Write \[ f(x)= a x^2+b x+c \]

OpenStudy (anonymous):

You have to find a , b c Solve \[ f(-2)=-14\\ f[0]=2\\ f(4) =-38 \] You have to solve 3 equations in three unknown. The 2nd equation gives c=2. So it is only two equations in two of them. Solve them and find a and b

OpenStudy (triciaal):

when i solved i got a = -3; b=-2 and c = 2 which would make f(x) -3x^2 -2x +2

OpenStudy (triciaal):

eqn 3 -38 = a(4)^2 + b(4) + 2 -10 = 4a + b eqn 1 -14 = a(-2)^2 + b(-2) +2 -16 = 4a -2b -8 = 2a - b add the 2 eqn -18 = 6a so a = -3 find b using eqn 1 -8 =2(-3) -b -8 + 6 = -2 a= -3; b = -2; c =2

OpenStudy (anonymous):

By the way b=2 and this is the final answer \[f(x)=-3 x^2+2 x+2\] Here is the graph of f and the points